Beg. inventory 10 units at $50
1st purchase 20 units at $60
2nd purchase 52 units at $65
3rd purchase 25 units at $63
4th purchase 37 units at $64


If Keisy sold107 units of inventory, calculate (show work) ending inventory and cost of goods sold using the LIFO method?
Answer #1

Ending inventory = Units available for sale - Units sold

= (10+20+52+25+37) - 107

= 144 - 107 = 37 units

Ending inventory

= (10*50) + (20*60) * (7*65)

= 500 + 1200 + 455

= 2155

Cost of Goods Sold

= (37*64) + (25*63) + (45*65)

= 2368 + 1575 + 2925

= 6868